# DualDynamics: Synergizing Implicit and Explicit Methods for Robust Irregular Time Series Analysis

### v3

@@ Introduction
Figure 5: Ablation study of different flow model configurations
We utilized two datasets, MuJoCo and Google, with different levels of observation for the forecasting task. Detailed explanations of each dataset and results are provided in Supplementary material.
MuJoCo.
We utilized the MuJoCo dataset (Todorov, Erez, and Tassa 2012), based on the Hopper configuration from the DeepMind control suite (Tassa et al. 2018). Following Jhin et al. (2023, 2024), we used 50 initial points to predict the next 10, introducing 30%, 50%, and 70% missing data scenarios. Table 5 illustrates the forecasting performance across diverse conditions, with our method consistently demonstrating minimal MSE.
Google.
Discussion
Performance versus network capacity.
Our analysis reveals that the performance improvements of DualDynamics are not merely a result of increased network capacity, but rather stem from the synergistic integration of implicit and explicit methods in our model architecture. To validate DualDynamics’ effectiveness, we conducted controlled experiments comparing it against models with similar parameter counts. Figures 3 and 4 illustrate the relationship between parameter count and MSE for our flow architecture versus conventional MLPs. Figure 3 shows that increasing MLP capacity yields limited improvement on the MuJoCo dataset, while DualDynamics maintains superior performance. Figure 4 demonstrates that with optimal flow configuration, DualDynamics consistently outperforms MLPs on the Google dataset across all model sizes.
This performance advantage persists even when adjusting baseline model complexity to match DualDynamics, indicating that our framework’s superiority stems from its architectural design rather than increased computational resources. These results highlight DualDynamics’ efficiency in capturing complex temporal dynamics in irregular time series data.
Ablation study of flow configurations.
Figure 5 presents a comprehensive overview of our ablation study, illustrating the performance of various flow configurations across different tasks, as detailed in the preceding sections. The results demonstrate a clear advantage of flow-based architectures over the conventional MLP approach, refer to the detailed results in the Supplementary material. These results suggest that incorporating implicit and explicit could lead to significant improvements in model performance and capabilities.
@@ Conclusion
We introduce DualDynamics, an innovative methodology that integrates implicit and explicit methods for modeling irregularly-sampled time series data. This sophisticated paradigm excels in capturing intricate temporal characteristics inherent in continuous-time processes, offering an enhanced framework for representation learning in time series analysis. By synergistically combining these approaches, our method strikes a balance between expressive power and computational efficiency, addressing key challenges in temporal modeling.
Our approach consistently surpasses existing benchmarks in classification, interpolation, and forecasting tasks, significantly elevating the precision and depth of understanding in temporal dynamics. These outstanding results highlight DualDynamics’ potential for diverse real-world applications. Its adaptability to various temporal modeling challenges positions it as a significant advancement in irregular time series analysis.
Ethic Statement
We commit to conducting our research with integrity, ensuring ethical practices and responsible use of technology in alignment with established academic and scientific standards.
@@ Acknowledgments
We thank the teams and individuals for their efforts in dataset preparation and curation for our research, especially the UEA & UCR repository for the numerous datasets we extensively analyzed.
@@
Appendix A Experiment details
We consider following experiments: (1) Interpolation for the missing data, (2) Classification of irregularly-sampled data, and (3) Forecasting with the partial observations. All experiments were performed using a server on Ubuntu 22.04 LTS, equipped with an Intel(R) Xeon(R) Gold 6242 CPU and multiple NVIDIA A100 40GB GPUs. The source code for our experiments can be accessed at https://github.com/yongkyung-oh/DualDynamics.
@@ Appendix A Experiment details
Neural CDEs provide a framework for constructing a continuous-time representation by formulating a controlled path $X(t)$, generated through natural cubic spline interpolation of observed data (Kidger et al. 2020). The Neural CDEs is expressed via the Riemann–Stieltjes integral:
@@
with the initial condition ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$ where $h:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{z}}$ is an affine function with parameter $\theta_{h}$, similar to Neural ODEs. Here, $f(t,{\bm{z}}(t);\theta_{f})$ represents the drift term, which models the deterministic part of the dynamics, while $g(t,{\bm{z}}(t);\theta_{g})$ is the diffusion term that captures the stochastic component. $W(t)$ denotes a standard Wiener process (Brownian motion). In Eq. (22), both $f$ and $g$ are modeled using neural networks parameterized by $\theta_{f}$ and $\theta_{g}$, respectively. Neural SDEs provide a framework for modeling complex, stochastic temporal dynamics, allowing for the representation of uncertainty in the latent state evolution.
@@
Training algorithm for DualDynamics framework
@@
Appendix B Detailed results for “Robust classification with the missing data”
@@
In our study, we conducted classification experiments across 18 diverse datasets, in three distinct domains, from the University of East Anglia (UEA) and the University of California Riverside (UCR) Time Series Classification Repository  [^4] http://www.timeseriesclassification.com/ (Bagnall et al. 2018) using the python library sktime (Löning et al. 2019). As outlined in Table 7, the datasets in the Sensor domain are univariate, whereas the others encompass a mix of univariate and multivariate data. Addressing the challenge posed by varying lengths of time series, the strategy of uniform scaling was employed to align all series to the dimension of the longest series (Keogh 2003; Yankov et al. 2007; Gao and Lin 2018; Tan et al. 2019). The partitioning of the dataset into training, validation, and testing subsets was executed in proportions of 0.70, 0.15, and 0.15 respectively. Subsequent to this partitioning, a random assortment of missing observations was introduced for each variable. Lastly, these modified variables are integrated into combined dataset.
Table 7: Description of datasets for the classification task
@@
Table 8: Performance and computation time using ‘BasicMotions’ dataset under regular (0% missingness) and irregular (30%, 50%, and 70% missingness) scenarios (Computation time is calculated without early-stopping)
@@
Table 8 presents the classification outcomes for both with and without missingness. While Neural CDE generally exhibits robust performance, its extended variants tend to require increased computation time and do not consistently perform well across a broad range of tasks compared to the benchmark performance. On the other hand, Neural Flow achieves rapid convergence with reduced performance. In contrast, our proposed method not only shows superior performance but also maintains a balanced computational demand.
Detailed results of ablation study
Table 9: Ablation study of the proposed method (Values in parentheses show average of 18 standard deviations)
@@
Variations in model architecture
Table 10: Comparative study of latent value selection
Configuration Regular 30% Missing 50% Missing 70% Missing
Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Proposed method
(Neural CDE)
Another consideration is the replacement of the flow model with more sophisticated architectures such as feedforward networks (Hornik, Stinchcombe, and White 1989), gated recurrent unit (Chung et al. 2014), temporal convolution networks (Lea et al. 2017), or transformer-based attention mechanisms (Vaswani et al. 2017). Table 11 summarizes the analysis of different model components in our framework. While these alternatives demonstrate strong performance on regularly-sampled time series, they often struggle with irregular sampling patterns and missing data. Compared to that, the flow model in our current architecture specifically addresses these challenges.
Table 11: Comparative study of component selection
Configuration Regular 30% Missing 50% Missing 70% Missing
Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Proposed method
(Neural CDE)
@@
Our investigation involved 18 datasets classified into three domains. We collate the outcomes across three domains: Motion & HAR, ECG & EEG, and Sensor datasets, in Tables 14, 14, and 14. We observe variability in the performance metrics due to the unique characteristics of each domain. Figure 7 shows the strong interdependence between the efficacy of our proposed approach and the distinctive attributes of the datasets under examination. In summary, the proposed methodology demonstrates superior performance relative to benchmark models.
(a) All domains
(b) Motion & HAR
(c) ECG & EEG
(d) Sensor
Figure 7: Comparison of performance across diverse domains and four missing rate scenarios
@@
Table 14: Average classification performance on 6 datasets in Sensor
@@
Appendix C Detailed results for “Interpolation for the missing data”
@@
The optimal hyperparameters for the proposed methods are remarked in bold within Table 16, encompassing observed data with 50%. We used the same hyperparameters to the remainder settings of 60%, 70%, 80%, and 90%.
@@ Appendix D Detailed results for “Classification of irregularly-sampled data”
We utilized the PhysioNet Sepsis dataset (Reyna et al. 2019) to investigate classification performance with irregularly-sampled time series. The dataset comprises 40,335 patient instances from ICU, characterized by 34 temporal variables such as heart rate, oxygen saturation, and body temperature.
Given that the PhysioNet dataset represents irregular time series, with only 10% of the data points timestamped for each patient, the study addresses this irregularity through two distinct approaches to time-series classification as suggested by Kidger et al. (2020): (i) classification utilizing observation intensity (OI) and (ii) classification without utilizing observation intensity (No OI). Observation intensity serves as an indicator of the severity of a patient’s condition, and when employed, each data point in the time series is supplemented with an index reflecting this intensity. Considering the dataset’s skewed distribution, performance is evaluated using the Area Under the Receiver Operating Characteristic Curve (AUROC) metric. Benchmark performance is established in Jhin et al. (2022).
Table 17: Ablation study on PhysioNet Sepsis dataset
@@
We followed experimental protocol suggested by Kidger et al. (2020), and its Github repository [^18] https://github.com/patrick-kidger/NeuralCDE. For the benchmark methods, we used the reported performance in Jhin et al. (2022). Table 17 presents a comprehensive ablation study comparing various flow configurations integrated with Neural CDEs. We train for $200$ epochs with a batch size of $1024$ and a learning rate of $0.001$. The best hyperparameters are highlighted with bold font in Table 21 and 21.
@@ Appendix E Detailed results for “Forecasting with the partial observations”
For the MuJoCo dataset, we followed experimental protocol suggested by in Jhin et al. (2023) and Jhin et al. (2024). We used reported performance of benchmark methods and source code from the Github repositories [^19] https://github.com/alflsowl12/LEAP. For the Google dataset, We followed experimental protocol and reported performance established by Jhin et al. (2022). Source code from its Github repository [^20] https://github.com/sheoyon-jhin/EXIT was utilized to conduct experiment. We conducted hyperparameter tuning for regular data and the same set of hyperparameters was applied to the scenarios with missingness.
@@
Table 18: Forecasting performance and memory usage on MuJoCo dataset
@@
Table 18 summarize the ablation study of the proposed method on MuJoCo dataset. In the proposed method, the training spanned $500$ epochs, with a batch size configured at $1024$ and a learning rate set to $0.001$. The hyperparameters that were identified as optimal are conspicuously presented in bold within Table 23.
@@
The Google stock data (Jhin et al. 2022) includes transaction volumes of Google, along with metrics such as high, low, open, close, and adjusted closing prices. This dataset spans from 2011 to 2021, and the objective is to utilize historical time-series data from the preceding 50 days to forecast the high, low, open, close, and adjusted closing prices for the next 10-day period. We followed the experimental protocol and reported performances of benchmark methods from Jhin et al. (2022). Similarly to the previous experiment, we excluded partial observations in data.
Table 19: Forecasting performance and memory usage on Google dataset
@@
Table 19 summarize the ablation study of the proposed method on Google dataset. In the proposed method, the training spanned $200$ epochs, with a batch size configured at $64$ and a learning rate set to $0.001$. Through grid search, the optimal hyperparameters are highlighted in bold within Table 23.
(Memory usage in MB, and the number of parameter in ×$\times 10^{3}$ unit.)
@@
Table 22: Results of hyperparameter tuning for MuJoCo Dataset
@@
Table 23: Results of hyperparameter tuning for Google Dataset

### v4

@@ Discussion
Figure 6: Ablation study of different flow model configurations versus conventional MLP
Ablation Study of Flow Model Configurations.
Figure 6 presents a comprehensive overview of our ablation study with four datasets, illustrating the performance of various flow configurations across different tasks, as detailed in the preceding sections. The results demonstrate a clear advantage of flow-based architectures over the conventional MLP approach, refer to the detailed results in the supplementary material. These results suggest that incorporating implicit and explicit could lead to significant improvements in model performance and capabilities.
@@ Conclusion
We introduce DualDynamics, an innovative methodology that integrates implicit and explicit methods for modeling irregularly-sampled time series data. This sophisticated paradigm excels in capturing intricate temporal characteristics inherent in continuous-time processes, offering an enhanced framework for representation learning in time series analysis. By synergistically combining these approaches, our method strikes a balance between expressive power and computational efficiency, addressing key challenges in temporal modeling for the irregular time series.
Our approach consistently surpasses existing benchmarks in classification, interpolation, and forecasting tasks, significantly elevating the precision and depth of understanding in temporal dynamics. These outstanding results highlight DualDynamics’ potential for diverse real-world applications. Its adaptability to various temporal modeling challenges positions it as a significant advancement in the irregular time series analysis in practices.
Ethics Statement
We commit to conducting our research with integrity, ensuring ethical practices and the responsible use of technology, fully aligning our efforts with established academic and scientific standards to promote trust and accountability.
@@ Acknowledgments
We thank the teams and individuals for their efforts in the real-world dataset preparation and curation for our research, especially the UEA & UCR repository for the numerous datasets that we extensively analyzed.
@@
Appendix A Experimental Details
We consider following experiments: (1) Robustness to Dataset Shift, (2) Classification of irregularly-sampled data, (3) Interpolation of Missing Data, and (4) Forecasting with the partial observations. All experiments were performed using a server on Ubuntu 22.04 LTS, equipped with an Intel(R) Xeon(R) Gold 6242 CPU and multiple NVIDIA A100 40GB GPUs. The source code for our experiments and datasets can be accessed at https://github.com/yongkyung-oh/DualDynamics.
@@ Appendix A Experimental Details
Neural CDEs provide a framework for constructing a continuous-time representation by formulating a control path $X(t)$, generated through natural cubic spline interpolation of observed data (Kidger et al. 2020). The Neural CDEs is expressed via the Riemann–Stieltjes integral:
@@
with the initial condition ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$ where $h:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{z}}$ is an affine function with parameter $\theta_{h}$, similar to Neural ODEs. Here, $f(t,{\bm{z}}(t);\theta_{f})$ represents the drift term, which models the deterministic part of the dynamics, while $g(t,{\bm{z}}(t);\theta_{g})$ is the diffusion term that captures the stochastic component. $W(t)$ denotes a standard Wiener process (or Brownian motion). In Eq. (22), both $f$ and $g$ are modeled using neural networks parameterized by $\theta_{f}$ and $\theta_{g}$, respectively. Neural SDEs provide a framework for modeling complex, stochastic temporal dynamics, allowing for the representation of uncertainty in the latent state evolution as continues-time perspective.
@@
Training Procedure
@@
Appendix B Detailed Results for “Robustness to Dataset Shift”
@@
We conducted classification experiments across 18 diverse datasets, in three distinct domains, from the University of East Anglia (UEA) and the University of California Riverside (UCR) Time Series Classification Repository  [^3] http://www.timeseriesclassification.com/ (Bagnall et al. 2018) using the python library sktime (Löning et al. 2019). As outlined in Table 7, the datasets in the Sensor domain are univariate, whereas the others encompass a mix of univariate and multivariate data. Addressing the challenge posed by varying lengths of time series, the strategy of uniform scaling was employed to align all series to the dimension of the longest series (Keogh 2003; Yankov et al. 2007; Gao and Lin 2018; Tan et al. 2019). The partitioning of the dataset into training, validation, and testing subsets was executed in proportions of 0.70, 0.15, and 0.15 respectively. Subsequent to this partitioning, a random assortment of missing observations was introduced for each variable. Lastly, these modified variables are integrated into combined dataset.
@@
Table 8 presents the classification outcomes for both with and without missingness. Computation times are measured with regular setting using a single GPU. While Neural CDE generally exhibits robust performance, its extended variants tend to require increased computation time and do not consistently perform well across a broad range of tasks compared to the benchmark performance. On the other hand, Neural Flow achieves rapid convergence with reduced performance. In contrast, our proposed method not only shows superior performance but also maintains a balanced computational demand.
@@
Table 8: Performance and computation time comparison using ‘BasicMotions’ dataset under regular (0% missingness) and irregular (30%, 50%, and 70% missingness) scenarios. Computation time is calculated without early-stopping.
Variations in model architecture
Another consideration is the replacement of the flow model with more sophisticated architectures such as feedforward networks (Hornik, Stinchcombe, and White 1989), gated recurrent unit (Chung et al. 2014), temporal convolution networks (Lea et al. 2017), or transformer-based attention mechanisms (Vaswani et al. 2017). Table 10 summarizes the analysis of different model components in our framework. While these alternatives demonstrate strong performance on regularly-sampled time series, they often struggle with irregular sampling patterns and missing data. Compared to that, the flow model in our current architecture specifically addresses these challenges. The flow model in our current architecture specifically addresses these challenges through its continuous and invertible nature, making it particularly well-suited for handling irregular time series data with missing observations.
Configuration Regular 30% Missing 50% Missing 70% Missing
Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Proposed method
(Neural CDE)
Table 9: Comparative study of latent value selection using ‘BasicMotions’ dataset
Configuration Regular 30% Missing 50% Missing 70% Missing
Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Proposed method
(Neural CDE)
[…]
– ResNet Flow 0.983 (0.037) 0.048 (0.053) 0.967 (0.075) 0.078 (0.121) 0.917 (0.102) 0.135 (0.121) 1.000 (0.000) 0.041 (0.035)
– GRU Flow 0.967 (0.046) 0.093 (0.073) 0.950 (0.075) 0.116 (0.107) 0.967 (0.046) 0.158 (0.096) 0.983 (0.037) 0.088 (0.057)
– Coupling Flow 1.000 (0.000) 0.007 (0.008) 1.000 (0.000) 0.021 (0.027) 0.983 (0.037) 0.044 (0.063) 0.983 (0.037) 0.069 (0.141)
Table 10: Comparative study of component selection using ‘BasicMotions’ dataset
@@
Table 11: Ablation study of the proposed method on 18 datasets under regular and three missing rates
(Values in parentheses show average of 18 standard deviations)
Detailed Results of Ablation Study
@@
Our investigation involved 18 datasets classified into three domains. We collate the outcomes across three domains: Motion & HAR, ECG & EEG, and Sensor datasets, in Tables 13, 13, and 14. We observe variability in the performance metrics due to the unique characteristics of each domain. Overall, the proposed methodology demonstrates superior performance relative to benchmark models, in the different domains and missing rate scenarios.
@@
Table 13: Average classification performance on 6 datasets in ECG & EEG
@@
Table 14: Average classification performance on 6 datasets in Sensor
Appendix C Detailed Results for “Interpolation of Missing Data”
@@
Hyperparameter optimization is conducted through a grid search, focusing on the number of layers $n_{l}\in\left\{1,2,3,4\right\}$ and hidden vector dimensions $n_{h}\in\left\{16,32,64,128\right\}$. The optimal hyperparameters for the proposed methods are remarked in bold within Table 16, encompassing observed data with 50%. We used the same hyperparameters to the remainder settings of 60%, 70%, 80%, and 90%.
@@ Appendix D Detailed Results for “Classification of Irregularly-Sampled Data”
We utilized the PhysioNet Sepsis dataset (Reyna et al. 2019) to investigate classification performance with irregularly-sampled time series. The dataset comprises 40,335 patient instances from ICU, characterized by 34 temporal variables such as heart rate, oxygen saturation, and body temperature. The objective is to detect the presence or absence of sepsis in each patient.
Given that the PhysioNet dataset represents irregular time series, with only 10% of the data points timestamped for each patient, the study addresses this irregularity through two distinct approaches to time series classification as suggested by Kidger et al. (2020): (i) classification utilizing observation intensity (OI) and (ii) classification without utilizing observation intensity (No OI). Observation intensity serves as an indicator of the severity of a patient’s condition, and when employed, each data point in the time series is supplemented with an index reflecting this intensity. Considering the dataset’s skewed distribution, performance is evaluated using the Area Under the Receiver Operating Characteristic Curve (AUROC) metric. Benchmark performance is established in Jhin et al. (2022).
@@
Table 17: Ablation study on PhysioNet Sepsis dataset
We followed experimental protocol suggested by Kidger et al. (2020), and its Github repository [^17] https://github.com/patrick-kidger/NeuralCDE. For the benchmark methods, we used the reported performance in Jhin et al. (2022). Table 17 presents a comprehensive ablation study comparing various flow configurations integrated with Neural CDEs. We train for $200$ epochs with a batch size of $1024$ and a learning rate of $0.001$. The hyperparameters are optimized by grid search in the number of layers in vector field $n_{l}\in\left\{1,2,3,4\right\}$ and hidden vector dimensions $n_{h}\in\left\{16,32,64,128\right\}$. The best hyperparameters are highlighted with bold font in Table 21 and 21. Here it is acknowledged that increasing $n_{h}$ magnify memory and complexity instead of $n_{l}$. However, a more complex model does not necessarily assure superior performance.
@@ Appendix E Detailed Results for “Forecasting with the Partial Observations”
We used two datasets for the forecasting with partial observations scenario. For the MuJoCo dataset, we followed experimental protocol suggested by in Jhin et al. (2023) and Jhin et al. (2024). We used reported performance of benchmark methods and source code from the Github repositories [^18] https://github.com/alflsowl12/LEAP. For the Google dataset, We followed experimental protocol and reported performance established by Jhin et al. (2022). Source code from its Github repository [^19] https://github.com/sheoyon-jhin/EXIT was utilized to conduct experiment. We conducted hyperparameter tuning for regular data and the same set of hyperparameters was applied to the scenarios entailing 30%, 50%, and 70% missing data.
@@
Table 18: Forecasting performance and memory usage on MuJoCo dataset
Table 18 summarize the ablation study of the proposed method on MuJoCo dataset. In the proposed method, the training spanned $500$ epochs, with a batch size configured at $1024$ and a learning rate set to $0.001$. The process of hyperparameter optimization was conducted through grid search, concentrating on the number of layers $n_{l}\in\left\{1,2,3,4\right\}$, and the dimensions of the hidden vector $n_{h}\in\left\{16,32,64,128\right\}$. The hyperparameters that were identified as optimal are conspicuously presented in bold within Table 23. The same set of hyperparameters was applied to the irregularly-sampled scenarios.
@@
The Google stock data (Jhin et al. 2022) includes transaction volumes of Google, along with metrics such as high, low, open, close, and adjusted closing prices. This dataset spans from 2011 to 2021, and the objective is to utilize historical time series data from the preceding 50 days to forecast the high, low, open, close, and adjusted closing prices for the next 10-day period. We followed the experimental protocol and reported performances of benchmark methods from Jhin et al. (2022). Similarly to the previous experiment, we excluded partial observations in data —30%, 50%, and 70%— in order to make more challenging situations.
@@
Table 19: Forecasting performance and memory usage on Google dataset
Table 19 summarize the ablation study of the proposed method on Google dataset. In the proposed method, the training spanned $200$ epochs, with a batch size configured at $64$ and a learning rate set to $0.001$. The process of hyperparameter optimization was conducted through grid search, concentrating on the number of layers $n_{l}\in\left\{1,2,3,4\right\}$, and the dimensions of the hidden vector $n_{h}\in\left\{16,32,64,128\right\}$. Through grid search, the optimal hyperparameters are highlighted in bold within Table 23. The same set of hyperparameters was applied to the irregularly-sampled scenarios.
@@
Table 22: Results of hyperparameter tuning for MuJoCo Dataset
@@
Table 23: Results of hyperparameter tuning for Google Dataset
(Memory usage in MB, and the number of parameter in ×$\times 10^{3}$ unit.)